# Bring Reason to Vision: Understanding Perception and Reasoning through Model Merging

### v1

@@ 7 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning Interpretability. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B More analysis for MathVista
Figure 10: Performance comparison of TIES and linear merging methods across MathVerse (Text-Dominant) and MathVision benchmarks. Baseline presents the result of LLaVA-Next-LLaMA3-8B. Linear shows the result of merging Dart-prop vector with hyperparameters of $(0.9,0.1)$. Performance reached by TIES merging with the same two task vectors are labeled as the hyperparameter pair used.
@@ Appendix C Different merging methods perform generally comparable
We are also interested in whether different merging methods affect the performance of reasoning ability transfer. To investigate this, we employ TIES (Yadav et al., 2023), a state-of-the-art merging method, to compare its performance with that of linear merging. Figure 10 shows that while TIES merging may demonstrate superior performance on certain benchmarks, such as MathVision and MathVerse (Text-Dominant), the overall average performance remains comparable to that of linear merging. Additionally, without the constraint that the sum of hyperparameters must be balanced across both task vectors, the hyperparameter search space expands significantly. As a result, TIES merging requires more hyperparameter tuning to achieve comparable performance.
MathVision MathVerse-Text-Dominant
Baseline 13.8 25.9
Linear 14.8 30.7
TIES(1, 0.4) 15.1 29.7
@@
Dare-TIES(0.8, 0.2) 17.8 22.7
Dare-Linear(0.8, 0.2) 17.4 21.7
Table 5: Performance comparison on MathVision and MathVerse-Text-Dominant datasets.
We also expand our experiments on Dare merging (Dare-TIES and Dare-Linear) and results shown in 5, another SOTA merging method to explore additional merging techniques. We adopt the same hyperparameter search strategy as TIES, parameterized by $(\alpha_{1},\alpha_{2})$, where $\alpha_{1}$ is tuned for the VLM and $\alpha_{2}$ for the Math LLM, to obtain the most comparable checkpoints on benchmarks emphasizing visual and textual reasoning. The results show that DARE performs comparably to linear merging and TIES, consistent with our finding that different merging methods yield similar performance, with none significantly outperforming simple averaging. This supports our choice to adopt the linear merging method. This finding is also consistent with the previous finding in Yadav et al. (2024) that different merging methods tend to exhibit similar behaviors at larger scales. Given this, we chose not to focus extensively on exploring alternative merging methods but instead to explore more about the interpretability and composition of the model’s internal abilities.
Appendix D More results on MM-Math
@@
Appendix E Additional results demonstrating generalization from logical to mathematical reasoning
Model Method MathVista MathVerse Benchmarks MV
All General Math Overall T−D T−L V−I V−D V−O
LLaVA Baseline 37.4 51.7 25.4 20.1 25.9 20.8 21.1 16.5 16.0 13.8
Appendix F Significance Test
To evaluate whether the improvements from merging are significant, we conduct a significance test. Merging general reasoning models like Mammoth2 and Magpie does not yield statistically significant enhancements, exhibiting a pattern of marginal accuracy gains, as shown in Table 2 of our paper. In contrast, merging math-focused LLMs, such as Dart-Prop, results in statistically significant improvements. This supports our conclusion that merging with math-related models offers the greatest advantage.
Model Method MathVista MathVerse MMStar DM
Math T-D T-L V-I V-D V-O Math Math
LLaVA +Dart-Uniform 0.15 0.00 0.00 0.00 0.02 0.46 0.56 0.21
+MAmmoTH-2 0.77 0.58 0.30 0.04 0.70 0.35 1.00 1.00
+Magpie-v0.3 0.67 1.00 0.06 0.07 0.84 0.62 0.56 0.96
+Dart-prop 0.06 0.00 0.01 0.00 0.01 0.02 0.05 0.00
Table 8: P-values across the models and datasets. Bolded values indicate $p<0.05$, which pass the significance test.

### v2

@@ 7 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning Interpretability. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix B More analysis for MathVista
Figure 10: Performance comparison of TIES and linear merging methods across MathVerse-Text-Dominant) and MathVision benchmarks. Baseline presents the result of LLaVA-Next-LLaMA3-8B. Linear shows the result of merging Dart-prop vector with hyperparameters of $(0.9,0.1)$. Performance reached by TIES merging with the same two task vectors are labeled as the hyperparameter pair used.
@@ Appendix C Different merging methods perform generally comparable
We are interested in whether different merging methods affect the performance of reasoning ability transfer. To investigate this, we employ TIES (Yadav et al., 2023), Dare merging (Dare-TIES and Dare-Linear) (Yu et al., 2024b), and Layer Swapping (Bandarkar et al., 2024), which are popular merging methods in practice, to compare their performance with that of linear merging. We adopt the hyperparameter search strategy as linear merging, parameterized by $(\alpha_{1},\alpha_{2})$, where $\alpha_{1}$ is tuned for the VLM vector and $\alpha_{2}$ for the Math LLM task vector, to obtain the most comparable checkpoints on benchmarks emphasizing visual and textual reasoning.
As shown in Figure 10, we present results of each configuration for TIES merging on both the visual- and text-dominant tasks, MathVision and MathVerse-Text-Dominant, accordingly. Compared to the performance yielded by linear merging, most combinations of TIES trade off gains in one domain against losses in the other. Additionally, the constraint of $\alpha_{1}+\alpha_{2}=1$ is commonly dropped in practice because of sparsification during TIES merging process. As a result, TIES merging requires more hyperparameter tuning to achieve comparable performance in a expanded searching space.
We also expand our experiments on the other methods and results shown in Table 5. The results performs comparably to TIES merging, consistent with our finding that different merging methods yield similar performance, with none significantly outperforming simple averaging. This supports our choice to adopt the linear merging method. This finding is also consistent with the previous finding in Yadav et al. (2024) that different merging methods tend to exhibit similar behaviors at larger scales. Given this, we chose not to focus extensively on exploring alternative merging methods but instead to explore more about the interpretability and composition of the model’s internal abilities.
Method Weights MathVision MathVerse-Text-Dominant
Baseline - 13.8 25.9
Linear (0.9, 0.1) 14.8 30.7
@@
Dare-TIES (1.0, 0.2) 17.8 22.7
Dare-Linear (1.2, 0.2) 17.4 21.7
Swap5Layers (0.9, 0.1) 15.1 29.6
Table 5: Performance comparison of different merging methods on MathVision and MathVerse-Text-Dominant datasets.
Appendix D Additional results demonstrating generalization from logical to mathematical reasoning
Model Method MathVista MathVerse Benchmarks MV
All General Math Overall T−D T−L V−I V−D V−O
LLaVA Baseline 37.4 51.7 25.4 20.1 25.9 20.8 21.1 16.5 16.0 13.8
Model Method MathVista MathVerse Benchmarks MMStar DM MV
All General Math Overall T−D T−L V−I V−D V−O All Math
Qwen2-VL Baseline 61.2 69.6 54.1 31.8 35.9 31.4 31.5 33.1 26.9 59.9 59.2 34.4 21.1
+Qwen2-Math 60.2 68.0 53.5 $\downarrow$0.6 31.9 $\uparrow$0.1 37.1 31.7 31.5 32.5 26.7 59.5 58.4 $\downarrow$0.8 35.0 $\uparrow$0.6 21.7 $\uparrow$0.6
Table 7: Performance (%) of Qwen2-VL before and after augmenting with Qwen2-Math on MathVista, MathVerse Benchmarks, MMStar, DM and MV. Arrows denote absolute change from the baseline.
Appendix E More results on Qwen2-VL
To assess the generality of this approach, we further extend our experiments to a stronger vision–language model, Qwen2-VL. In Table 7, we present the results of merging Qwen2-VL-7B-Instruct with Qwen2-Math-7B. The merged model shows significant improvements on MathVerse, DynaMath and MathVision, but a performance drop on MathVista and MMStar. This uneven pattern mirrors our observations with Idefics and InternVL, suggesting that VLMs already pretrained on mathematics-focused text corpora—as is common in many state-of-the-art models today—derive less benefit from integration with specialized reasoning modules (Laurençon et al., 2024).
Appendix F More results on MM-Math
@@
Appendix G Significance Test
To evaluate whether the improvements from merging are significant, we conduct a significance test and mark in Table LABEL:tab:_sig_test. Merging general reasoning models like Mammoth2 and Magpie does not yield statistically significant enhancements, exhibiting a pattern of marginal accuracy gains, as shown in Table 2 of our paper. In contrast, merging math-focused LLMs, such as Dart-Prop, results in statistically significant improvements. This supports our conclusion that merging with math-related models offers the greatest advantage.
Model Method MathVista MathVerse Benchmarks MMStar DM
Math Overall T-D T-L V-I V-D V-O Math
LLaVA Baseline 25.4 20.1 25.9 20.8 21.1 16.5 16.0 30.0 22.7
+Dart-Uniform 28.3 $\uparrow$2.9 23.6 $\uparrow$3.5 32.0 25.6 25.4 19.3 17.4 31.6 $\uparrow$1.2 24.5 $\uparrow$1.8
+MAmmoTH-2 25.7 $\uparrow$0.3 20.6 $\uparrow$0.5 26.0 22.0 22.1 16.4 16.1 30.0 $\uparrow$0.0 22.5 $\downarrow$0.2
+Magpie-v0.3 25.9 $\uparrow$0.5 20.7 $\uparrow$0.6 26.8 22.2 22.6 16.2 15.5 30.8 $\uparrow$0.8 22.7 $\uparrow$0.0
+Dart-prop 28.9 $\uparrow$3.5 23.7 $\uparrow$3.6 30.7 24.8 25.5 19.5 17.4 33.6 $\uparrow$3.6 24.5 $\uparrow$1.8
Table 9: Significant test results across the models and datasets. marks statistically significant improvements ($p<0.05$).